# Frontier AI’s Impact on the Cybersecurity Landscape

### v3

@@ Abstract
The impact of frontier AI in cybersecurity is rapidly increasing. In this paper, we comprehensively analyze this trend through three distinct lenses: a quantitative benchmark analysis, a literature review, and an expert survey. We find that while AI is already widely used in attacks, its application in defense remains limited, especially in remediation and deployment. Aligned with these analyses, experts expect AI to continue favoring attackers over defenders, though the gap will gradually narrow. These findings underscore the urgent need to mitigate frontier AI’s risks while closely monitoring emerging capabilities. We provide concrete calls-to-action regarding: the construction of new cybersecurity benchmarks, the development of AI agents for defense, the design of provably secure AI agents, the improvement of pre-deployment security testing and transparency, and the strengthening of user-oriented education and defenses. Our paper summary and blog are available at https://rdi.berkeley.edu/frontier-ai-impact-on-cybersecurity/.
@@ 1 Introduction
The security community is increasingly using frontier AI, [^1] We use frontier AI to represent foundation models like large language models (LLMs) and multi-modal models, as well as AI agents [66]—systems that combine foundation models with symbolic software components. e.g., using LLMs for fuzzing [156] and patching [276]. Researchers have also developed foundation models for security tasks such as analyzing networking traffic, binary code, and blockchain transactions, achieving better performance than traditional ML in threat detection [91, 77, 262, 179]. More recently, AI agents from the DARPA AIxCC competition [54] discovered and patched zero-day vulnerabilities [93]. Despite these defensive advances, frontier AI introduces significant risks, particularly its potential for misuse by attackers—ranging from the automation of cyber attacks [121, 236, 71, 223] to attacks against AI systems [237, 22], e.g, backdoors [164], jailbreaking [39, 40], and prompt injection [146].
Given this dual nature, a key question arises: Will frontier AI provide more benefits to attackers or defenders, and how might it reshape the cybersecurity landscape? While several recent papers examine LLMs in cybersecurity [73, 98, 81, 261, 193, 274, 280, 108, 205, 196], these studies do not address these questions: (1) How can we evaluate the current state of frontier AI in attacks and defenses? (2) How might the offense-defense balance change with frontier AI? and (3) What actions need to be taken to ensure AI benefits defense, increasing the security of systems? Answering these questions is especially critical to prepare for increasing access to frontier AI, amid the unprecedentedly rapid pace of AI evolvement (Figure 1[2, 202].
Figure 1: Performance evolvement of frontier AI models over time. The coding-related tasks are emphasized in bold text.
@@
In this paper, we present a comprehensive analysis through benchmark analysis, a literature review, and an expert survey. First, we define the marginal risks of frontier AI to cybersecurity [122] and propose a risk taxonomy and risk assessment method (Section 2). We then apply this taxonomy to assess frontier AI’s current impacts on both attacks and defenses quantitatively and qualitatively (Sections 3 and 4). Our findings show that frontier AI has a much stronger influence on the attack side—particularly in human-targeted attacks—while its role in defense remains limited. In particular, remediation and deployment currently show no measurable impact. Through the expert survey, we also find that experts believe that frontier AI will benefit attackers more, even though the gap is expected to narrow over time; within the next two years, the balance is expected to favor attackers by $31.7\%,$ within five years by $12.8\%,$ and within ten years by just $0.46\%$ (Section 5). Our position aligns with this view, which we support through analysis of the equivalence class problem and the fundamental asymmetry between attackers and defenders (Section 6). Finally, we offer actionable recommendations to guide the community as follows (Section 7):
@@ 2 Methodology
Based on the concept of marginal risk [122]—the new/unique risks beyond the existing risks—we define the marginal risks of frontier AI in cybersecurity as novel cyber threats either introduced or amplified by frontier AI through intentional misuse beyond traditional attack methods. For attacks against systems, we categorize attacks into seven steps based on the MITRE framework [161] and the cyber kill chain framework [151] (Figure 2); within each step, we define subcategories by activity and application. Similarly, we classify attacks against humans based on their goals and methods, as well as defenses for systems. We acknowledge that covering all risks and attack vectors is challenging; therefore, we limit our scope to cyberattacks against systems and humans (e.g, we do not discuss cryptographic attack vectors). With the taxonomy, we study frontier AI’s impact as follows:
Quantitative benchmarking. Because the evaluation of frontier AI’s cybersecurity capabilities is still at an early stage, available benchmarks remain limited. To collect benchmarks, we begin with several early and representative cybersecurity benchmarks (e.g., Cybench [272]), selected based on our domain expertise. We expand this set by adding other benchmarks cited within those works. We compile $34$ cybersecurity benchmarks and categorize them based on our taxonomy. We also identify the limitations of current benchmarks and propose calls-to-action to improve benchmarking in Section 7.
@@ 3 Quantitative Analysis: Benchmarks
1 Recon. 1 AutoPenBench [80] Limited method and scenario coverage; limited agent scaffolds
@@
Mal. C 2 RedCode [88], CySecBench [235] Limited metric for malware creation,
Vul. E 7 CyberSecEval [236], BaxBench [232], BountyBench [270], CyberGym [242] limited program language coverage
@@
Ini. A & Del. 2 AutoPenBench [80], SecCodePLT [260] Limited attack category and type coverage,
Ex. & Inst. 3 CyBench [272], SecCodePLT [260] limited benchmarks on real-world complex systems,
@@
Pen. test 1 AutoPenBench [80] Limited attacks, target systems, metrics,
Vul. D 12 PrimeVul [63], SVEN [100] FuzzBench [159], OSS-Fuzz [200] limited complexity and diversity, low-quality labels
2 Att. D 4 IDS2018 [204], Drebin [16], BODMAS [259] Low-quality data and labels, lack OOD/hidden tests
@@
PoC & RT 2 CRUXEval [85], CyberGym [242] No benchmark for root cause
4 Rem. Dev. 5 SWE-bench [118], SecCodePLT [260], BountyBench [270] Limited benchmarks for vuln. in different languages
@@ 3.1 Offensive Benchmarks
Attack Step 1: Reconnaissance. Early benchmarks focus on question answering [225, 147, 236, 124]. AutoPenBench [80] enables attack generation and execution but covers limited methods and scenarios. Results show that simple AI agents can handle reconnaissance like discovering potential victims on remote networks or conducting surveillance. However, agents cannot identify target-specific vulnerable services due to limited service information from standard tools (e.g., Nmap) and a lack of deeper interactions with target systems.
Attack Step 2: Weaponization. For malware creation, RedCode [88] tests LLMs’ capabilities in generating malware in Python and uses VirusTotal as the judge. CySecBench [235] tests the safety alignment of LLMs and shows that Claude 3.5 Sonnet is more resilient than GPT-4o and Gemini-1.5. Recent works design benchmarks for insecure code generation (e.g., CyberSecEval [236], SecCodePLT [260], BaxBench [232], SecRepoBench [62]). They craft specific coding tasks that may encounter known CWEs and evaluate whether LLMs generate vulnerable code when finishing these tasks, showing that both SOTA commercial and open-source models can generate vulnerable code, with stronger models being more vulnerable due to their enhanced task understanding and execution. BountyBench [270] and CVE-Bench [282] evaluate agent performance on vulnerability exploitation, showing that agents using Claude 3.7 Sonnet successfully exploit $67.5\%$ of tasks on BountyBench, while agents using GPT-4o achieve a $13\%$ success rate on CVE-Bench. CyberGym [242] evaluates agents’ capabilities in PoC generation on real-world software projects, identifying $15$ zero-day vulnerabilities. All existing benchmarks focus on C/C++ with limited coverage of other languages.
@@ 3.2 Defensive Benchmarks
Defense Step 1: Proactive testing. Benchmarking penetration testing is related to attack benchmarking, which remains at an early stage with AutoPenBench [80]. Existing vulnerability detection benchmarks include static benchmarks with labeled functions [176, 63, 228, 35, 226, 234, 92, 172] and dynamic benchmarks with fuzzing instrumentation and execution environment [200, 159]. For static benchmarks, most existing benchmarks remain at the individual function level, e.g., PrimeVul [63] and SVEN [100]. These datasets have low-quality labels due to missing context (functions vulnerable only under conditions specified by other functions in the projection). SecCodePLT [260] includes a project-level vulnerability detection dataset for C/C++ and Java, showing the limited performance of existing models and agents in these languages. For dynamic benchmarks, OSS-Fuzz [200] is the largest continuous fuzzing platform for critical open-source projects. FuzzBench [159] is a smaller but less noisy benchmark widely used in research papers to test novel fuzzing methods, including AI-facilitated fuzzers [208, 275].
@@
Defense Step 2: Attack detection. Many benchmarks exist for network intrusion detection and malware detection [204, 16, 259]. Recent research [15] highlights these benchmarks’ limitations in data quality (duplicated data, shortcuts) and label accuracy. Transformers show very high accuracy on these benchmarks [91, 140]. Besides attack detections, CTIBench [7] designs tasks for LLMs to pinpoint the specific CWEs and attack techniques for given attack instances.
@@
Defense Step 3: Triage & forensic. CRUXEval [85] benchmarks PoC generation for vulnerable C/C++ codes with overflow bugs, showing that GPT-4o achieves 75% pass@1 success rate. CyberGym [242] offers more challenging tasks for agents, including PoC generation on real-world codebases with vulnerability descriptions, showing that OpenHands with Claude 4 Sonnet achieves a 17.9% success rate.
@@
Defense Steps 4&5: Remediation development and deployment. Given that AI is not yet widely involved in these steps, there are not many benchmarks. The most popular benchmark is SWE-bench [118], which extracts issues from multiple GitHub Python projects. SOTA multi-agent systems with GPT-5 and Claude 4 Sonnet can resolve more than $70\%$ issues in the SWE-bench-verified benchmark. Recent works extend the SWE-bench to multi-modal settings [256] and the Java language [267]. SecCodePLT [260] (Python, C/C++, Java) and BountyBench [270] (C, Python, JavaScript/TypeScript) include benchmarks of security patches. Both benchmarks use unit tests to test functionality. SecCodePLT tests security via fuzzing through PoC, and BountyBench tests whether the patch prevents human-written exploits. Top agents (e.g., OpenAI Codex CLI o4-mini) can achieve high results on BountyBench (>90%), but remain ineffective on SecCodePLT (<30%).
@@
1: Reconnaissance Scan the env. or harvest info. to identify targets [221, 56, 189][160] LLMs in reconnaissance include active scanning, victim information gathering, and open-source database search.
Step 2: Weaponization Create malware and exploits and encapsulate them as weapons [173, 205, 149, 135, 194, 63] [143, 236, 255, 242, 71, 270] [72, 175, 223, 272, 203] [20, 230, 160, 93, 115, 245] Research papers show LLMs can generate functional malware with high evasion rates and AI agents aid in vulnerability identification and exploitation, although large-scale real-world attacks remain limited to simpler attacks like DDoS and SQL injection.
Step 3: Delivery Transmit attacks through e-mail, USB, web [107, 238, 104, 181, 57, 53] [75, 169, 115, 160] Initial access are demonstrated in the real world, with AI assisting in password guessing and bypassing defenses such as 2FA.
Step 4: Exploitation Exploit vulnerabilities in the victim [52] AI can write simple exploits but cannot craft exploit chains.
Step 5: Installation Install the malware on the victim’s system [23][83, 199] AI is used for malware installation but not persistence in the real world.
Step 6: Command & control Establish a remote command channel [95, 52, 227, 125, 96, 199, 11] [23, 43][119, 198, 171] Privilege escalation through exploit chain generation and command&control through automated domain generation are demonstrated in research. Credential access is used in the real world.
Step 7: Action on objectives Achieve their ultimate objectives (e.g. data theft, sabotage) through the C&C channel [260, 250, 239, 229, 211] [20, 109, 230, 115, 75, 169, 48] [269, 78, 30, 152, 79] Real-world AI-enhanced attacks increase across various systems (Web, mobile, cloud), with malicious purposes including malware deployment, business logic abuse, and credential theft, causing significant financial losses and data breaches.
Attacks against humans Exploit people often through deception, coercion, or manipulation [32, 70, 25, 101, 68, 14, 13, 82, 47] [182][36, 51, 213, 261, 197, 27] [153, 165, 210, 218, 106, 192, 18] Frontier AI significantly escalates attacks against humans (e.g., studies showing increases in social engineering and voice phishing since ChatGPT’s adoption).
@@
Step 1: Proactive testing Probe systems before an attack to uncover vulnerabilities [123, 94, 250, 224, 45, 105, 116] [56, 281, 252, 63, 4, 163, 194] [149, 135, 206, 183, 65, 220, 130] [205, 254, 168, 156, 58, 273, 275] [112, 136, 207, 155][269, 89] Research explores LLMs for proactive testing, including automated penetration testing and vulnerability detection and AI-enhanced fuzzing. Real-world demonstrations exist but lack evidence of large-scale adoption.
Step 2: Attack detection Identify malicious behavior as it happens and raise real-time alerts [240, 247, 34, 233, 127, 144, 140] [6, 91, 262, 77, 216, 266][33] LLMs remove manual feature engineering, reduce reliance on labeled datasets, and generalize better to out-of-distribution cases, with real-world use in malware and intrusion detection.
Step 3: Triage & forensic Reproduce the vulnerabilities and identify root causes [241, 99, 242, 166, 90, 162, 50] [191, 253, 184, 209, 42, 179, 248] [117, 251, 110][49] For PoC generation, recent works mainly explore LLM-assisted fuzzing. AI agents are developed for root cause analysis. Foundation models also show superior performance in binary analysis.
Step 4: Rem. dev. Create patches or other countermeasures [133, 263, 126, 3, 157, 214, 41, 177] [276, 28, 139, 167, 265, 132][93] Research demonstrates AI can automatically generate security vulnerability patches, but real-world applications remain limited
@@
Defense for humans Protect people from manipulation or exploitation in attacks [74, 268, 60, 31, 187, 5, 12] [61, 142, 277, 174, 141, 278, 87] [131, 158] Even though frontier AI shows promise for enhancing defenses against human-targeted attacks, defensive techniques struggle to keep pace with sophisticated attacks.

### v4

@@ Abstract
The impact of frontier AI (i.e., AI agents and foundation models) in cybersecurity is rapidly increasing. In this paper, we comprehensively analyze this trend through multiple aspects: quantitative benchmarks, qualitative literature review, empirical evaluation, and expert survey. Our analyses consistently show that AI’s capabilities and applications in attacks have exceeded those on the defensive side. Our empirical evaluation of widely used agent systems on cybersecurity benchmarks highlights that current AI agents struggle with flexible workflow planning and using domain-specific tools for complex security analysis—capabilities particularly critical for defensive applications. Our expert survey of AI and security researchers and practitioners indicates a prevailing view that AI will continue to benefit attackers over defenders, though the gap is expected to narrow over time. These results show the urgent need to evaluate and mitigate frontier AI’s risks, steering it towards benefiting cyber defenses. Responding to this need, we provide concrete calls to action regarding: the construction of new cybersecurity benchmarks, the development of AI agents for defense, the design of provably secure AI agents, the improvement of pre-deployment security testing and transparency, and the strengthening of user-oriented education and defenses. Our paper summary and blog are available at https://rdi.berkeley.edu/frontier-ai-impact-on-cybersecurity/.
@@ 1 Introduction
The security community is increasingly using frontier AI, [^1] We use frontier AI to represent foundation models like large language models (LLMs) and multi-modal models, as well as AI agents [durante2024agent]—systems that combine foundation models with symbolic software components. e.g., using LLMs for fuzzing [meng2024large] and patching [zhang2024evaluating]. Researchers have also developed foundation models for security tasks such as analyzing networking traffic, binary code, and blockchain transactions, achieving better performance than traditional ML in threat detection [guthula2023netfound, gai2023blockchain, yu2024blockfound, pei2020xda]. More recently, AI agents from the DARPA AIxCC competition [darpa] discovered and patched zero-day vulnerabilities [SQLite3]. Despite these defensive advances, frontier AI introduces significant risks, particularly its potential for misuse by attackers—ranging from the automation of cyber attacks [kang2024exploiting, wan2024cyberseceval, fang2024llm, tann2023using] to attacks against AI systems [wang2023decodingtrust, barrett2023identifying], e.g, backdoors [nie2024trojfm], jailbreaking [chen2024llm, chen2024rl], and prompt injection [liu2023prompt].
Given this dual nature, a key question arises: Will frontier AI provide more benefits to attackers or defenders, and how might it reshape the cybersecurity landscape? While several recent papers examine LLMs in cybersecurity [ferrag2024generative, hassanin2024comprehensive, girhepuje2024survey, yao2024survey, sai2024generative, zhang2025llms, zhou2024large, hou2024large, sheng2025large, schroer2025sok], these studies do not address these questions: (1) How can we evaluate the current state of frontier AI in attacks and defenses? (2) How might the offense-defense balance change with frontier AI? and (3) What actions need to be taken to ensure AI benefits defense, increasing the security of systems? Answering these questions is especially critical to prepare for increasing access to frontier AI, amid the unprecedentedly rapid pace of AI evolvement (Figure 1[aisafetyreport, shah2025approach].
Figure 1: Software engineering performance evolvement of frontier AI over time.
@@
Our work presents a comprehensive analysis through benchmark analysis, literature review, empirical evaluation, and expert survey. First, we describe a marginal risk framework [kapoor2024societal] for analyzing AI’s impact in cybersecurity (Section 2). We then apply this taxonomy to assess frontier AI’s current impacts on both attacks and defenses quantitatively and qualitatively (Sections 3 and 4). Our analysis indicates that AI has been utilized more in attacks—particularly in human-targeted attacks—than in defense, enabling greater automation and facilitation to attackers.
We also empirically evaluate the state-of-the-art (SOTA) coding agent OpenHands [wang2024openhands] on the popular attack and defense benchmarks (Section 5). The results support our earlier conclusion and further uncover the limitations of current agents in architecture and workflow planning. Specifically, the agent is not equipped with domain-specific tools, especially those beneficial for defenses, which prevents it from conducting security analyses on large codebases. As a result, the agent performs poorly in difficult tasks, such as exploit development, proof-of-concept (PoC) generation, and root cause analysis (if not in stack traces).
Through the expert survey, we also find that experts believe that frontier AI will benefit attackers more, even though the gap is expected to narrow over time; within the next two years, the balance is expected to favor attackers by $31.7\%,$ within five years by $12.8\%,$ and within ten years by just $0.46\%$ (Section 6). Our position aligns with this view, which we support through analysis of the equivalence class problem and the fundamental asymmetry between attackers and defenders (Section 7).
Finally, we offer actionable recommendations to guide the community as follows (Section 8):
@@ 2 Methodology
Marginal risk framework and scope. Based on the concept of marginal risk [kapoor2024societal]—the new/unique risks beyond the existing risks—we define the marginal risks of frontier AI in cybersecurity as novel cyber threats either introduced or amplified by frontier AI through intentional misuse beyond traditional attack methods. For attacks against systems, we categorize attacks into seven steps based on the MITRE framework [mitre_attack_2024] and the cyber kill chain framework [kill_chain_2024] (Figure 2); within each step, we define subcategories by activity and application. Similarly, we classify attacks against humans based on their goals and methods, as well as defenses for systems. We acknowledge that covering all risks and attack vectors is challenging; therefore, we limit our scope to cyberattacks against systems and humans (e.g, we do not discuss cryptographic attack vectors). With the taxonomy, we study frontier AI’s impact as follows:
Quantitative benchmarking. Because the evaluation of frontier AI’s cybersecurity capabilities is still at an early stage, available benchmarks remain limited. To collect benchmarks, we begin with several early and representative cybersecurity benchmarks (e.g., Cybench [zhang2024cybench]), selected based on our domain expertise. We expand this set by adding other benchmarks cited within those works. We compile $34$ cybersecurity benchmarks and categorize them based on our taxonomy. We also identify the limitations of current benchmarks and propose calls-to-action to improve benchmarking in Section 8.
@@
Empirical Evaluation. To examine the limited defensive capabilities of current agents relative to their offensive capabilities, we evaluate the offensive and defensive capabilities of the SOTA agent, OpenHands [wang2024openhands], using several cybersecurity benchmarks. For offensive capability, we employ AutoPenBench [gioacchini2024autopenbench], which includes end-to-end attack scenarios. For defensive capability, we use CyberGym [wang2025cybergym] for PoC generation and PatchAgent [yu2025patchagent] for patch generation. Detailed experimental setups and designs are described in Section 5.
@@ 3 Quantitative Analysis: Benchmarks
1 Recon. 1 AutoPenBench [gioacchini2024autopenbench] Limited method and scenario coverage; limited agent scaffolds
@@
Mal. C 2 RedCode [guo2024redcode], CySecBench [wahreus2025cysecbench] Limited metric for malware creation,
Vul. E 7 CyberSecEval [wan2024cyberseceval], BaxBench [vero2025baxbench], BountyBench [zhang2025bountybench], CyberGym [wang2025cybergym] limited program language coverage
@@
Ini. A & Del. 2 AutoPenBench [gioacchini2024autopenbench], SecCodePLT [yang2024seccodeplt] Limited attack category and type coverage,
Ex. & Inst. 3 CyBench [zhang2024cybench], SecCodePLT [yang2024seccodeplt] limited benchmarks on real-world complex systems,
@@
Pen. test 1 AutoPenBench [gioacchini2024autopenbench] Limited attacks, target systems, metrics,
Vul. D 12 PrimeVul [ding2024vulnerability], SVEN [he2023large] FuzzBench [metzman2021fuzzbench], OSS-Fuzz [serebryany2017oss] limited complexity and diversity, low-quality labels
2 Att. D 4 IDS2018 [sharafaldin2018toward], Drebin [arp2014drebin], BODMAS [yang2021bodmas] Low-quality data and labels, lack OOD/hidden tests
@@
PoC & RT 2 CRUXEval [gu2024cruxeval], CyberGym [wang2025cybergym] No benchmark for root cause
4 Rem. Dev. 5 SWE-bench [jimenez2023swe], SecCodePLT [yang2024seccodeplt], BountyBench [zhang2025bountybench] Limited benchmarks for vuln. in different languages
@@ 3.1 Offensive Benchmarks
Attack Step 1: Reconnaissance. Early benchmarks focus on question answering [tihanyi2024cybermetric, liu2024cyberbench, wan2024cyberseceval, kouremetis2025occult]. AutoPenBench [gioacchini2024autopenbench] enables attack generation and execution but covers limited methods and scenarios. Results show that simple AI agents can handle reconnaissance like discovering potential victims on remote networks or conducting surveillance. However, agents cannot identify target-specific vulnerable services due to limited service information from standard tools (e.g., Nmap) and a lack of deeper interactions with target systems.
Attack Step 2: Weaponization. For malware creation, RedCode [guo2024redcode] tests LLMs’ capabilities in generating malware in Python and uses VirusTotal as the judge. CySecBench [wahreus2025cysecbench] tests the safety alignment of LLMs and shows that Claude 3.5 Sonnet is more resilient than GPT-4o and Gemini-1.5. Recent works design benchmarks for insecure code generation (e.g., CyberSecEval [wan2024cyberseceval], SecCodePLT [yang2024seccodeplt], BaxBench [vero2025baxbench], SecRepoBench [dilgren2025secrepobench]). They craft specific coding tasks that may encounter known CWEs and evaluate whether LLMs generate vulnerable code when finishing these tasks, showing that both SOTA commercial and open-source models can generate vulnerable code, with stronger models being more vulnerable due to their enhanced task understanding and execution. BountyBench [zhang2025bountybench] and CVE-Bench [zhu2025cvebench] evaluate agent performance on vulnerability exploitation, showing that agents using Claude 3.7 Sonnet successfully exploit $67.5\%$ of tasks on BountyBench, while agents using GPT-4o achieve a $13\%$ success rate on CVE-Bench. CyberGym [wang2025cybergym] evaluates agents’ capabilities in PoC generation on real-world software projects, identifying $15$ zero-day vulnerabilities. All existing benchmarks focus on C/C++ with limited coverage of other languages.
@@ 3.2 Defensive Benchmarks
Defense Step 1: Proactive testing. Benchmarking penetration testing is related to attack benchmarking, which remains at an early stage with AutoPenBench [gioacchini2024autopenbench]. Existing vulnerability detection benchmarks include static benchmarks with labeled functions [pearce2022asleep, ding2024vulnerability, ullah2024llms, chauvin2024eyeballvul, tony2023llmseceval, vulhub, Hackthebox, owasp] and dynamic benchmarks with fuzzing instrumentation and execution environment [serebryany2017oss, metzman2021fuzzbench]. For static benchmarks, most existing benchmarks remain at the individual function level, e.g., PrimeVul [ding2024vulnerability] and SVEN [he2023large]. These datasets have low-quality labels due to missing context (functions vulnerable only under conditions specified by other functions in the projection). SecCodePLT [yang2024seccodeplt] includes a project-level vulnerability detection dataset for C/C++ and Java, showing the limited performance of existing models and agents in these languages. For dynamic benchmarks, OSS-Fuzz [serebryany2017oss] is the largest continuous fuzzing platform for critical open-source projects. FuzzBench [metzman2021fuzzbench] is a smaller but less noisy benchmark widely used in research papers to test novel fuzzing methods, including AI-facilitated fuzzers [shi2024harnessing, zhang2025g2fuzz].
@@
Defense Step 2: Attack detection. Many benchmarks exist for network intrusion detection and malware detection [sharafaldin2018toward, arp2014drebin, yang2021bodmas]. Recent research [arp2022and] highlights these benchmarks’ limitations in data quality (duplicated data, shortcuts) and label accuracy. Transformers show very high accuracy on these benchmarks [guthula2023netfound, lin2022bert]. Besides attack detections, CTIBench [alam2024ctibench] designs tasks for LLMs to pinpoint the specific CWEs and attack techniques for given attack instances.
@@
Defense Step 3: Triage & forensic. CRUXEval [gu2024cruxeval] benchmarks PoC generation for vulnerable C/C++ codes with overflow bugs, showing that GPT-4o achieves 75% pass@1 success rate. CyberGym [wang2025cybergym] offers more challenging tasks for agents, including PoC generation on real-world codebases with vulnerability descriptions, showing that OpenHands with Claude 4 Sonnet achieves a 17.9% success rate.
@@
Defense Steps 4&5: Remediation development and deployment. Given that AI is not yet widely involved in these steps, there are not many benchmarks. The most popular benchmark is SWE-bench [jimenez2023swe], which extracts issues from multiple GitHub Python projects. SOTA multi-agent systems with GPT-5 and Claude 4 Sonnet can resolve more than $70\%$ issues in the SWE-bench-verified benchmark. Recent works extend the SWE-bench to multi-modal settings [yang2024swe] and the Java language [zan2024swe]. SecCodePLT [yang2024seccodeplt] (Python, C/C++, Java) and BountyBench [zhang2025bountybench] (C, Python, JavaScript/TypeScript) include benchmarks of security patches. Both benchmarks use unit tests to test functionality. SecCodePLT tests security via fuzzing through PoC, and BountyBench tests whether the patch prevents human-written exploits. Top agents (e.g., OpenAI Codex CLI o4-mini) can achieve high results on BountyBench (>90%), but remain ineffective on SecCodePLT (<30%).
@@
1: Reconnaissance Scan the env. or harvest info. to identify targets [sun2024gptscan, deng2024pentestgpt, nvidia_cve_analysis][Microsoft] LLMs in reconnaissance include active scanning, victim information gathering, and open-source database search.
Step 2: Weaponization Create malware and exploits and encapsulate them as weapons [pa2023attacker, sheng2025large, lu2024grace, li2024llm, sakaoglu2023kartal, ding2024vulnerability] [liumake, wan2024cyberseceval, yang2023exploitgen, wang2025cybergym, fang2024llm, zhang2025bountybench] [fang2024teams, patil2024leveraging, tann2023using, zhang2024cybench, shao2024empirical] [ddos, a10ddos2024, Microsoft, SQLite3, imperva_news, google_bigsleep] Research papers show LLMs can generate functional malware with high evasion rates and AI agents aid in vulnerability identification and exploitation, although large-scale real-world attacks remain limited to simpler attacks like DDoS and SQL injection.
Step 3: Delivery Transmit attacks through e-mail, USB, web [hitaj2019passgan, wang2023password, securityhero2023, plesner2024breaking, deng2024oedipusllmenchancedreasoningcaptcha, debicha2023advbot] [gmail_account_news, openai_report_disrupting, imperva_news, Microsoft] Initial access are demonstrated in the real world, with AI assisting in password guessing and bypassing defenses such as 2FA.
Step 4: Exploitation Exploit vulnerabilities in the victim [de2024chainreactor] AI can write simple exploits but cannot craft exploit chains.
Step 5: Installation Install the malware on the victim’s system [beckerich2023ratgpt][google_adv_misuse, deeplocker] AI is used for malware installation but not persistence in the real world.
Step 6: Command & control Establish a remote command channel [happe2024llmshackersautonomouslinux, de2024chainreactor, tulla2025alfa, kujanpaa2021automating, happe2023evaluating, deeplocker, anderson2016deepdga] [beckerich2023ratgpt, chung2019availability][obviam-password-cracking, securityhero-password-cracking, cepa-ai-cyberattacks] Privilege escalation through exploit chain generation and command&control through automated domain generation are demonstrated in research. Credential access is used in the real world.
Step 7: Action on objectives Achieve their ultimate objectives (e.g. data theft, sabotage) through the C&C channel [yang2024seccodeplt, xu2024autoattacker, wang2024sands, usman2024generative, singer2025feasibility] [ddos, hp_news, a10ddos2024, imperva_news, gmail_account_news, openai_report_disrupting, crowdstrike] [zero, gartner_survey, bugcrowd_inside, ai-retail, ai-retail2] Real-world AI-enhanced attacks increase across various systems (Web, mobile, cloud), with malicious purposes including malware deployment, business logic abuse, and credential theft, causing significant financial losses and data breaches.
Attacks against humans Exploit people often through deception, coercion, or manipulation [phishing, eze2024analysis, bethany2024large, heiding2024evaluating, lawenforcggment, spearphishing, anthropic_persuasiveness, goldstein2024persuasive, costello2024durably] [potter2024hidden][hongkong, darktrace, voice_phishing, yao2024survey, fraudgpt, blauth2022artificial] [theft, openai_report_disrupting_25feb, fakereview, sullivan2024ai, cyberbullying, psy_control, suicide] Frontier AI significantly escalates attacks against humans (e.g., studies showing increases in social engineering and voice phishing since ChatGPT’s adoption).
@@
Step 1: Proactive testing Probe systems before an attack to uncover vulnerabilities [kong2025vulnbot, happe2023getting, xu2024autoattacker, temara1maximizing, constantin2023llms, hilario2024generative, isozaki2024towards] [deng2024pentestgpt, zhou2019devign, xu2017neural, ding2024vulnerability, akuthota2023vulnerability, nana2024deep, sakaoglu2023kartal] [lu2024grace, li2024llm, shestov2024finetuning, purba2023software, du2024vul, sun2024llm4vuln, lekssays2025llmxCPG] [sheng2025large, yang2023kernelgpt, oliinyk2024fuzzing, meng2024large, deng2023large, zhang2024llamafuzz, zhang2025g2fuzz] [huang2024large, li2025iris, shi2024bandfuzz, meng2024largehybrid][zero, repoaudit2025] Research explores LLMs for proactive testing, including automated penetration testing and vulnerability detection and AI-enhanced fuzzing. Real-world demonstrations exist but lack evidence of large-scale adoption.
Step 2: Attack detection Identify malicious behavior as it happens and raise real-time alerts [wang2017adversary, wu2023grim, caviglione2020tight, vinayakumar2019robust, kwon2019survey, liu2025analyzing, lin2022bert] [al2024exploring, guthula2023netfound, yu2024blockfound, gai2023blockchain, stein2024towards, zahan2024shifting][microsoft_ire] LLMs remove manual feature engineering, reduce reliance on labeled datasets, and generalize better to out-of-distribution cases, with real-world use in malware and intrusion detection.
Step 3: Triage & forensic Reproduce the vulnerabilities and identify root causes [wang2024python, jingxuan2021learch, wang2025cybergym, nitin2025faultline, guo2019deepvsa, mu2019renn, cui2018rept] [roy2024exploring, yang2024large, rafi2024enhancing, shin2015recognizing, chua2017neural, pei2020xda, xie2024resym] [jiang2025beyond, xu2025unleashing, hu2024degpt][charlotte] For PoC generation, recent works mainly explore LLM-assisted fuzzing. AI agents are developed for root cause analysis. Foundation models also show superior performance in binary analysis.
Step 4: Rem. dev. Create patches or other countermeasures [li2025sok, yu2024security, kulsum2024case, ahmed2023better, meng2024empirical, sobania2023analysis, chen2024large, pearce2023examining] [zhang2024evaluating, bouzenia2024repairagent, lin2024vulnerabilities, nong2025appatch, yu2025patchagent, li2025patchpilot][SQLite3] Research demonstrates AI can automatically generate security vulnerability patches, but real-world applications remain limited
@@
Defense for humans Protect people from manipulation or exploitation in attacks [ferrara2023social, zanke2023ai, dhieb2020secure, cao2024phishagent, rana2022deepfake, al2023gpt, anirudh2023multilingual] [dhiman2024gbert, liu2024preventing, zhang2024remark, pan2024markllm, liu2024survey, zhao2024sok, gunn2024undetectable] [nvdia, granny] Even though frontier AI shows promise for enhancing defenses against human-targeted attacks, defensive techniques struggle to keep pace with sophisticated attacks.